OpenStudy (anonymous):

if the triangle ABC, E is the midpoint of AC and D is the midpoint of CB. If DF is Parallel to BE, Find the length of AB?

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

so than D is the midpoint of BC and CD is equal 5 so this mean that BC = 10 writing property of sides you get BE/CB = DF/CD so BE/10 = 3/5 so 5BE = 30 so BE = 6 you can writing property of sides so CF/DF = CE/BE so 4/3 = CE/6 so 24=3CE so CE=8 than CE=8 then AC= 16 using property of sides you can writing AC/BC = AB/BE so 16/10 = AB/6 than 10AB = 96 result AB = 96/10 = 9,6 AB=9,6 hope this will help you